OpenStudy (anonymous):

If f(x) = log3 (x + 1), what is f^−1(2)?

hartnn (hartnn):

can you find f^(-1) (x) first ?

hartnn (hartnn):

y=log_3 (x+1) isolate x from here

OpenStudy (anonymous):

so would it be log3*1 = x

hartnn (hartnn):

what ? do u know \(if, \log_ab=c \implies a^c=b \) ?

OpenStudy (anonymous):

no

hartnn (hartnn):

now you know, so from that what about \(\log_3(x+1)=y \implies ???\)

OpenStudy (anonymous):

c^y = ( x + 1 ) .... i just took the formula you gave me and plugged in

hartnn (hartnn):

you mean \(\huge 3^y=x+1\) instead of c ?

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

and f^(-1) (2) means the value of y is given to be 2 so you just need 'x' from \(\huge 3^2=x+1\) got this ? any doubts till here ?

OpenStudy (anonymous):

1 8 10 27 these are my answer choices

hartnn (hartnn):

what is 3^2 =... ?

OpenStudy (anonymous):

9

hartnn (hartnn):

yes, so can u find x from 9=x+1 ??

OpenStudy (anonymous):

x=8

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

Given the function f(x) = 0.3(4)^x, what is the value of f^−1(6)? will i do the same for this

hartnn (hartnn):

6 = 0.3 4^x yes! same :)

OpenStudy (anonymous):

do i multiply 0.3 and 4

hartnn (hartnn):

nopes, 6/0.3 = 4^x

OpenStudy (anonymous):

i keep getting 1.03

hartnn (hartnn):

what are the choices ?

OpenStudy (anonymous):

0.360 1.242 2.161 2.312

hartnn (hartnn):

with 4^x = 6/0.3 using a calculator, i get 2.161

OpenStudy (anonymous):

what do i put in first

hartnn (hartnn):

put where ?

OpenStudy (anonymous):

the answer was correct thank you

OpenStudy (anonymous):

& i have the TI 84plus calculator ..i'm still trying to work it lol

hartnn (hartnn):

its ok :) and welcome ^_^

OpenStudy (anonymous):

would you like t help me on another assignment

hartnn (hartnn):

ask your questions in a new post, so that even if i am unavailable, others can help you :)

OpenStudy (anonymous):

it's up